# Key Words.: cosmology: distance scale – gravitational lensing: strong – Galaxy: kinematics and dynamics – Galaxies: elliptical and lenticular, cD1 Introduction

### v1

@@ 4.3.1 Prior on stellar temperature and metallicity
We test the effects of restricting the templates based on stellar age and metallicity (e.g., age between 1 and 12 Gyrs, and metallicity above 0.1 solar), in order to avoid potential non-physical templates. We have performed this test on SSP libraries as the age of the stars in stellar libraries is not available from the catalogs in the literature. The results of the tests shown in Figure 9 demonstrate that in some cases, non-physical templates (e.g., templates with ages of 15 Gyrs for the XSL SSP library) can be chosen while using the ‘full’ libraries. However, the effect on the inferred stellar velocity dispersion is minimal, well below our target of 1%, with very few exceptions for the XSL models. We thus conclude that it is not necessary to impose a priori cuts on age and metallicity, even though one should be careful in interpreting the results.
@@ 4.4.1 Stellar libraries
The dominant source of potential systematic errors is the choice of templates. Figures 10 to 14 show heat maps representing the average difference in stellar velocity dispersion between any pair of combination of stellar templates. For the MUSE, KCWI, and SDSS samples, the difference averaged over the galaxy sample is reported, while for JWST-NIRSpec, the average difference over 21 spaxels is reported.
It is clear that, it is possible to achieve sub-percent agreement among the measurements from the clean stellar libraries, irrespective of the galaxy sample or instrument being considered. The full stellar libraries can be significantly off, highlighting the importance of using clean libraries. This is particularly striking in the case of JWST-NIRSpec, where the differences between clean and full stellar libraries can be 10-15%.
@@ 4.4.2 Single stellar populations models
As shown in Figures 10 to 14, the stellar velocity dispersions based on the SSP models can differ by a few percent from those inferred based on stellar libraries. The difference is more pronounced for the KCWI and SDSS samples than for the MUSE sample. While for the former the SSP libraries give consistently higher stellar velocity dispersions by 2-4%, the differences are smaller for the MUSE sample. In order to investigate the origin of this discrepancy, we considered a subset of the MUSE sample matched in redshift and stellar velocity dispersion to the KCWI sample. For example, if we restrict the MUSE sample to and 220 km s-1 km s-1, then the stellar vs. SSP ratios get higher and closer to what we found for KCWI: Indo-US vs EMILES SSP: from -1.70.3% (MUSE-full) to -2.90.8% (MUSE-restricted), similar to -3.40.4% obtained for KCWI. The corresponding numbers for MILES vs. EMILES-SSP are: -2.90.3%, -3.60.8%, and %. For XSL vs. EMILES SSP they are:-2.4 0.3%, -3.90.8%, and -4.40.4%. In all cases, it seems that the SSP “bias” is larger when selecting the higher galaxies at lower redshifts. In general, those will have older ages and higher metallicities than their higher-z and lower- counterparts, suggesting deficiencies of the SSP libraries in this part of the parameter space. Our sample size is too small to carry out a detailed investigation of the bias between stellar and SSP libraries as a function of galaxy parameters. Hopefully, future studies will address this.
For the time being, we recommend avoiding the SSP models for high precision/accuracy measurements for a number of reasons. First, the SSPs are based on the assumption that spectra are single stellar populations, which is not true in general. Even a linear combination of a few of them may not have sufficient flexibility to describe the star formation history of the galaxy to the level we require. Indeed, as shown above, the SSPs lead to unphysical solutions in some cases, which raises a red flag. Second, the smaller number of SSP templates compared to stellar templates gives less flexibility to reproduce the diversity in metallicity and elemental abundance. Third, the SSPs are a combination of theoretical and empirical spectra, and we think it is more prudent to stick with empirical spectra for high-precision measurements, given the uncertainties in theoretical atmospheres.
In conclusion, SSPs are fine if one is interested in uncertainties at the 2-3% level, but we find that they are not appropriate for sub-percent accuracy.
@@ 4.5 Wavelength range
The wavelength range used for the fit is often dictated by practical considerations such as instrumental setup, wavelength coverage of the templates, and the need to avoid features from contaminants (3.3). Within these constraints, it is best to choose a wavelength region that contains deep absorption features such as CaHK, the G-band, and the Ca II Triplet. As we discussed above, the choice of polynomial orders to describe multiplicative and additive continuum terms depends on the wavelength region chosen (see Section 4.2). Fortunately, our extensive tests show that once appropriate polynomial orders are selected and clean stellar libraries are employed, the specific choice of the spectral region is a subdominant contribution to the systematic error budget. Small changes in wavelength coverage (i.e., shifts by 100 Å or so) are absolutely negligible for the quality of the data presented here. Even larger changes make a subdominant contribution to the error budget (see also D’Ago et al. 2023, for a similar result).
@@
We can take advantage of the SDSS spectra available from the SDSS database for the 14 galaxies with KCWI spectra to verify the effects of changing significantly the wavelength range. The wavelength ranges adopted vary substantially (3600-4500 Å for KCWI and 3600-5350 Å for SDSS), with the SDSS including important redder features such as the MgB triplet and Fe5270. We cannot go redder with SDSS because the XSL library is not sufficiently well behaved across the dichroic for our target level of precision and accuracy.
@@
In order to ensure a fair comparison, following Knabel et al. (2024), we convolve the KCWI data with a kernel to match the SDSS seeing and then measure the stellar velocity dispersion within a circular aperture of diameter . The SDSS error dominates the uncertainty, since the SDSS SNR is substantially lower than KCWI. For the Indo-US and XSL clean libraries, the stellar velocity dispersions agree within one , while for MILES, the discrepancy is 2.5. Given the uncertainties, we see no significant evidence for a difference, although the uncertainty is relatively large. This comparison should be taken with caution and should be repeated with better-quality data, ideally with a larger sample. Knabel et al. (2024) discusses this comparison in more detail. The main conclusion is that the SDSS data for these specific galaxies are not of sufficient quality to measure stellar velocity dispersions with precision and accuracy better than 5%, confirming the initial assessment by (Bolton et al. 2008).
@@
Shajib et al. (2025, in prep) will present a comparison between kinematics obtained with JWST in the Ca II Triplet region and with KCWI in the blue part of the spectrum. Knabel et al. (2024) presents more comparisons between SDSS and KCWI.
@@
In our analysis, we will not discuss the wavelength region much further, although in our suggested recipe, we include a reminder to check for the effects of the wavelength range.
@@
In summary, we consider the following rest-frame spectral ranges. For MUSE and KCWI, we adopt the range 3600-4500 Å , where the quality of the template libraries sets the blue cutoff, and the red cutoff is imposed by the redshift of the targets and the instrumental setup. For JWST, we use the range 8303-8838 Å for reasons discussed in Section 3.3. For SDSS, we consider 3600-5350 , taking advantage of the redder coverage to increase the information content of the spectra and avoiding complications from the UVB-VIS dichroic in the XSL templates.
@@ 5 A recipe for accurate stellar velocity dispersions
This section describes a recipe to quantify and minimize residual systematic errors and covariance. The recipe consists of a number of steps that will be summarized at the end of the section. At the core of the recipe is a well-established statistical method to marginalize over systematic uncertainties based on information contained in the data themselves. The reader not interested in the statistical details can skip directly to the final end product described in Section 6 and in Table 2, where we summarize the results obtained with our method, and those that one would obtain by the simplest approach – a straight average of the fits obtained with the three clean stellar libraries.
In practice, we consider the possible choice of template as a nuisance parameter and marginalize over it weighted by the probability of the model given the data. However, since the data and model parameters vary, to obtain a fair marginalization we cannot use a simple likelihood ratio, but we need to use a more general metric. Following, e.g., Birrer et al. (2019), we assign the statistical weight based on the Bayesian information criterion (BIC), defined as (Schwarz 1978):
where is the number of free parameters in the model, represents the number of data points (or pixels), and stands for the maximum value of the likelihood function, corresponding to the model.
Note that for a sample of galaxies, one has to construct the BIC for the whole sample since the first term of the right-hand side of Equation 4 is non-linear in . In other words, if there are galaxies in the sample with data points each, and they are each fit with degrees of freedom resulting in likelihood the equation becomes:
@@ 5.1 Counting free parameters
The definition of the number of free parameters in the BIC is non-trivial and requires some discussion. The parameter is the number of spectral points used in the fit, after masking. The number of free model parameters is given by , where the M is typically 2 and accounts for the stellar velocity dispersion and line-of-sight velocity (but it could be more if there are multiple components, see, e.g., Shajib et al. 2025), is the number of corrective polynomial orders (both additive and multiplicative), and is the number of “significant templates,” i.e., the number of nonzero templates, or in practice the number of templates that contribute above a certain threshold (in our case 1%) to the fit. The threshold is needed for two reasons. First, stellar libraries can include a large number of spectra that are not relevant to fit the data (e.g., B stars), and those should not be counted in order to obtain a fair representation of the effective freedom of the fit.
Second, from a numerical standpoint, the rationale for omitting the weights of zero or nearly-zero templates when estimating free parameters is that when a template weight reaches the boundary at the optimal fit, the positivity constraint in the template fit behaves like an equality constraint. This effectively eliminates one free parameter. During the constrained quadratic template optimization, pPXF identifies the optimal weights by aiming to minimize the to machine precision. This process can result in some very small weights, which, while formally necessary to optimize , make no significant contribution to the fit from a physical perspective.
Intuitively, the threshold is related to the noise level of the galaxy spectra. For our data quality SNR 30–160 Å-1 (i.e., 1-3% per Å), we find that 1% is the appropriate threshold using the following procedure. We first carry out a pPXF fit with all the templates in the library, and then we repeat the fit by keeping the spectra that contribute only above a certain threshold to the initial fit. As expected, we find that for a high threshold (10-20%), the BIC is large because there are not enough templates to obtain a good fit. Conversely, below %, the does not improve enough to compensate for the increased number of degrees of freedom in the BIC. This threshold depends on the quality of the data, approximately in a manner inversely proportional to the SNR. We recommend following our procedure to set the threshold for each specific dataset.
@@ 5.2 Computing the weights
The probability of a model given the data is formulated as:
Hence, the relative probability between two fits, and , given the data, can be written in terms of the fits’ corresponding BIC values:
However, we only have a noisy estimate of the BIC, with the noise given by statistical noise and sampling noise. Thus, we need to define the evidence ratio function in terms of a variable sampling the “true” difference between a fit’s corresponding BIC value and the lowest BIC among all possible fit configurations as discussed above, BICmin,k:
To find a model’s associated statistical weight, , we marginalize over , with the probability of given the measured described by a Gaussian, , centered on BICk with a variance of .
The standard deviation, , represents the uncertainty in the calculation and is dominated by numerical effects and population sampling. There could be, for example, one galaxy with unusual abundance ratios that is the best fit by a lucky template and skews the BIC. In contrast to the strong lensing cases described by Birrer et al. (2019) and Shajib et al. (2022), convergence is not an issue for the kinematic fits described here. The BIC uncertainty can then simply be estimated by bootstrap resampling the sample of galaxies. Importantly, there is covariance between the BIC calculated using a pair of stellar libraries: some galaxies may be better fit by more than one template library than others, and therefore, one needs to compute the scatter in BIC by bootstrap resampling the sample of galaxies, rather than computing the scatter independently for each library. Typically, we find for the samples under consideration here, although clearly, the BIC and its uncertainty depend on sample size.
@@
When the number of template libraries is small (as in the case considered here), it is prudent to adopt Bessel’s correction to get the unbiased estimators (see Appendix A for a derivation)
For a sample of galaxies or multiple spatial bins within a galaxy, in addition to measuring the average the mean , one can estimate the the covariance matrix as:
@@
For a small number of stellar libraries, again, we recommend Bessel’s corrected estimator, which reads

### v2

@@ 4.3.1 Prior on stellar temperature and metallicity
We tested the effects of restricting the templates based on stellar age and metallicity (e.g., age between 1 and 12 Gyrs, and metallicity above 0.1 solar), in order to avoid potential non-physical templates. We performed this test on SSP libraries as the age of the stars in stellar libraries is not available from the catalogs in the literature. The results of the tests shown in Figure 9 demonstrate that in some cases, non-physical templates (e.g., templates with ages of 15 Gyrs for the XSL SSP library) can be chosen while using the ”full” libraries. However, the effect on the inferred stellar velocity dispersion is minimal, well below our target of 1%, with very few exceptions for the XSL models. We thus conclude that it is not necessary to impose a priori cuts on age and metallicity, even though one should be cautious when interpreting the results.
@@ 4.4.1 Stellar libraries
The dominant source of potential systematic errors is the choice of templates. Figures 10 and 14 to 16 show heat maps representing the average difference in stellar velocity dispersion between any pair of combination of stellar templates. For the MUSE, KCWI, and SDSS samples, the difference averaged over the galaxy sample is reported, while for JWST-NIRSpec, the average difference over 21 spaxels is reported.
It is clear that it is possible to achieve sub-percent agreement among the measurements from the clean stellar libraries, irrespective of the galaxy sample or instrument being considered. The full stellar libraries can be significantly off, highlighting the importance of using clean libraries. This is particularly striking in the case of JWST-NIRSpec, where the differences between clean and full stellar libraries can be 10-15%.
@@ 4.4.2 Single stellar populations models
As shown in Figures 10 and 14 to 16, the stellar velocity dispersions based on the SSP models can differ by a few percent from those inferred based on stellar libraries. The difference is more pronounced for the KCWI and SDSS samples than for the MUSE sample. While for the former the SSP libraries give consistently higher stellar velocity dispersions by 2-4%, the differences are smaller for the MUSE sample. In order to investigate the origin of this discrepancy, we considered a subset of the MUSE sample matched in redshift and stellar velocity dispersion to the KCWI sample. For example, if we restrict the MUSE sample to and 220 km s-1 km s-1, then the stellar vs. SSP ratios get higher and closer to what we found for KCWI: Indo-US vs EMILES SSP: from -1.70.3% (MUSE-full) to -2.90.8% (MUSE-restricted), similar to -3.40.4% obtained for KCWI. The corresponding numbers for MILES vs. EMILES-SSP are: -2.90.3%, -3.60.8%, and %. For XSL vs. EMILES SSP they are:-2.4 0.3%, -3.90.8%, and -4.40.4%. In all cases, it seems that the SSP “bias” is larger when selecting the higher galaxies at lower redshifts. In general, those will have older ages and higher metallicities than their higher-z and lower- counterparts, suggesting deficiencies of the SSP libraries in this part of the parameter space. Our sample size is too small to carry out a detailed investigation of the bias between stellar and SSP libraries as a function of galaxy parameters. Hopefully, future studies will address this.
For the time being, we recommend avoiding the SSP models for high precision/accuracy measurements for a number of reasons. First, the SSPs are based on the assumption that spectra are single stellar populations, which is not true in general. Even a linear combination of a few of them might not have sufficient flexibility to describe the star formation history of the galaxy to the level we require. Indeed, as shown above, the SSPs lead to unphysical solutions in some cases, which raises a red flag. Second, the smaller number of SSP templates compared to stellar templates gives less flexibility to reproduce the diversity in metallicity and elemental abundance. Third, the SSPs are a combination of theoretical and empirical spectra, and we think it is more prudent to stick with empirical spectra for high-precision measurements, given the uncertainties in theoretical atmospheres.
In conclusion, SSPs are fine if one is interested in obtaining uncertainties at the 2-3% level. However, we find that they are not appropriate for the sub-percent accuracy required for cosmology.
@@ 4.5 Wavelength range
The wavelength range used for the fit is often dictated by practical considerations such as instrumental setup, wavelength coverage of the templates, and the need to avoid features from contaminants (3.3). Within these constraints, it is best to choose a wavelength region that contains deep absorption features such as CaHK, the G-band, and the Ca II Triplet. As we discussed above, the choice of polynomial orders to describe multiplicative and additive continuum terms depends on the wavelength region chosen (see Section 4.2). Fortunately, our extensive tests show that once appropriate polynomial orders are selected and clean stellar libraries are employed, the specific choice of the spectral region is a subdominant contribution to the systematic error budget. Small changes in wavelength coverage (i.e., shifts by 100 Å or so) are absolutely negligible for the quality of the data presented here. Even larger changes make a subdominant contribution to the error budget (see also DAGO23, for a similar result).
@@
We took advantage of the SDSS spectra available from the SDSS database for the 14 galaxies with KCWI spectra to verify the effects of changing significantly the wavelength range. The wavelength ranges we adopted vary substantially (3600-4500 Å for KCWI and 3600-5350 Å for SDSS), with the SDSS including important redder features such as the MgB triplet and Fe5270. We could not go redder with SDSS because the XSL library is not sufficiently well behaved across the dichroic for our target level of precision and accuracy. To ensure a fair comparison, following Knabel24, we convolved the KCWI data with a kernel to match the SDSS seeing and then measured the stellar velocity dispersion within a circular aperture of diameter . The SDSS error dominates the uncertainty, since the SDSS S/N is substantially lower than KCWI. The results are shown in Figure 11. For the Indo-US and XSL clean libraries, the stellar velocity dispersions agree within one , while for MILES, the discrepancy is 2.5. Given the uncertainties, we see no significant evidence to support a difference, although the uncertainty is relatively large.
@@
This comparison should be taken with caution and should be repeated with better quality data, ideally with a larger sample. Knabel24 discussed this comparison in more detail. The main conclusion is that the SDSS data for these specific galaxies are not of sufficient quality to measure stellar velocity dispersions with precision and accuracy better than 5%, confirming the initial assessment by (Bolton08).
@@
Shajib25 presented a comparison between kinematics obtained with JWST in the Ca II Triplet region and with KCWI in the blue part of the spectrum. Knabel24 presented further comparisons between SDSS and KCWI.
@@
In our analysis, we chose not to discuss the wavelength region much further. However, in our suggested recipe, we include a reminder to check for the effects of the wavelength range.
@@
In summary, we consider the following rest-frame spectral ranges. For MUSE and KCWI, we adopted the range 3600-4500 Å , where the quality of the template libraries sets the blue cutoff and the red cutoff is imposed by the redshift of the targets and the instrumental setup. For JWST, we used the range 8303-8838 Å for reasons discussed in Section 3.3. For SDSS, we considered 3600-5350 Å, taking advantage of the redder coverage to increase the information content of the spectra and avoiding complications from the UVB-VIS dichroic in the XSL templates.
@@ 5 A recipe for accurate stellar velocity dispersions
This section describes a recipe to quantify and minimize residual systematic errors and covariance. The recipe consists of a number of steps that will be summarized at the end of the section. At the core of the recipe is a well-established statistical method to marginalize over systematic uncertainties based on information contained in the data themselves. This section is focused on statistical details, followed by a presentation of the final end product described in Section 6 and in Table 2, where we summarize the results obtained with our method. We also include results we would obtain by the simplest approach, namely, a straight average of the fits obtained with the three clean stellar libraries.
In practice, we consider the possible choice of template library as a nuisance parameter and marginalize over it weighted by the probability of the model given the data. However, since the data and model parameters vary, to obtain a fair marginalization we cannot use a simple likelihood ratio, but we need to use a more general metric. Following, for instance, Birrer19b, we assigned the statistical weight, , based on the Bayesian information criterion (BIC), defined as (Schwarz1978):
where is the number of free parameters in the model, represents the number of data points (or pixels) used in the fit after masking, and stands for the maximum value of the likelihood function, corresponding to the model.
When the template libraries have been properly cleaned, the resulting likelihoods from fits to a single spectrum will generally not distinguish a preferred library. Instead, we determine the combined BIC for a sample of galaxies or for a set of spatially binned spectra for a single galaxy observed with IFU spectroscopy. Note that for a sample of galaxies or set of bins, we have to construct the BIC for the whole sample since the first term of the right-hand side of Equation 4 is non-linear in . In other words, if there are galaxies in the sample with data points each, and they are each fit with degrees of freedom resulting in likelihood the equation becomes
@@
All of the methods we present are equally valid for the case of a sample of galaxies and for the case of a set of spatial bins measured across a single galaxy. In the following text, where we refer to a sample of galaxies, we could also refer to the spatial bins of a single galaxy. In general, for spatially resolved samples like KCWI and MUSE, we would prefer to treat each object individually. However, none of the objects in our KCWI and MUSE samples have sufficient S/N across a large enough number of spatial bins to be able to strongly distinguish a BIC-preferred template library given the uncertainties in the BIC comparisons. For this reason, and the reasons outlined in Section 3, we integrate the datacubes to achieve the highest possible S/N for each object in the sample. Therefore, in the results described in Section 6, the preference for a specific template library is driven by the properties of the sample as a whole, including properties specific to the instrument used to obtain the data. However, we do not consider any of these BIC preferences to be universally applicable to data obtained with that instrument, nor to the same sample of galaxies observed with another instrument. Complete studies of spatially resolved datasets should also be assessed at the level of spatial bins. The procedures will be exactly the same as outlined in this section.
@@ 5.1 Counting free parameters
The definition of the number of free parameters in the BIC is non-trivial and requires some discussion. The number of free model parameters, , is given by , where the M is typically 2 and accounts for the stellar velocity dispersion and line-of-sight velocity (but it could be more if there are multiple components, see, e.g., Shajib25), is the number of corrective polynomial orders (both additive and multiplicative), and is the number of “significant templates,” namely, the number of nonzero templates, or in practice the number of templates that contribute above a certain threshold (in our case 1%) to the fit. The threshold is needed for two reasons. First, stellar libraries can include a large number of spectra that are not relevant to fit the data (e.g., B stars) and those should not be counted in order to obtain a fair representation of the effective freedom of the fit.
Second, from a numerical standpoint, the rationale for omitting the weights of zero (or nearly zero) templates when estimating free parameters is that when a template weight reaches the boundary at the optimal fit, the positivity constraint in the template fit behaves like an equality constraint. This effectively eliminates one free parameter. During the constrained quadratic template optimization, pPXF identifies the optimal weights by aiming to minimize the to machine precision. This process can result in some very small weights, which, while formally necessary to optimize , make no significant contribution to the fit from a physical perspective.
The count of effective degrees of freedom attributable to the templates used in the fit is therefore paramount to the proper comparison of model BICs. Intuitively, the threshold is related to the noise level of the galaxy spectra. For our data quality S/N of 30–160 Å-1 (i.e., 1-3% per Å), we find that 1% is the appropriate threshold using the following procedure. We first carry out a pPXF fit with all the templates in the library, and then we repeat the fit by keeping the spectra that contribute only above a certain threshold to the initial fit. As expected, we find that for a high threshold (10-20%), the BIC is large because there are not enough templates to obtain a good fit. Conversely, below %, the does not improve enough to compensate for the increased number of degrees of freedom in the BIC. This threshold depends on the quality of the data, approximately in a manner inversely proportional to the S/N. We recommend following our procedure to set the threshold for each specific dataset. Thomas_Lipka22 offered an alternative approach to counting the effective number of free parameters.
@@ 5.2 Computing the weights
In the following formulas, we use subscripts to label the models (e.g., template libraries) that are being compared, while refers to the sample of galaxy spectra or spatial bins. The probability of a model, , given the data is formulated as
Hence, the relative probability between two fits, and , given the data, can be written in terms of the fits’ corresponding BIC values,
However, we only have a noisy estimate of the BIC, with the noise given by statistical noise and sampling noise. Thus, we need to define the evidence ratio function in terms of a dummy variable sampling the “true” difference between a fit’s corresponding BIC value and the lowest BIC among all possible fit configurations as discussed above, BICmin,k,
To find a model’s associated statistical weight, , we marginalized over , with the probability of given the measured described by a Gaussian, , centered on BICk with a variance of . size. In the absence of uncertainty, would be strictly positive. However, when convolving for the effect of the uncertainty, the width of the convolution kernel (set by ) might extend some part of the kernel’s tail to , which we allow.
The standard deviation, , represents the uncertainty in the calculation and is dominated by numerical effects and population sampling. There could be, for example, one galaxy with unusual abundance ratios that is the best fit by a lucky template and skews the BIC. In contrast to the strong lensing cases described by Birrer19b and Shajib22b, convergence is not an issue for the kinematic fits described here. The BIC uncertainty can then simply be estimated by bootstrap resampling the sample of galaxies. Importantly, there is covariance between the BIC calculated using a pair of stellar libraries: some galaxies might be better fit by more than one template library than others, and therefore, we need to compute the scatter in BIC by bootstrap resampling the sample of galaxies, rather than computing the scatter independently for each library. Typically, we find for the samples under consideration here, although it is clear that the BIC and its uncertainty depend on sample size.
From the bootstrapped distributions for the SDSS sample, we test the null hypothesis that they are derived from the Gaussian distribution . Using for the three clean libraries (Indo-US, MILES, and XSL), we found Kolmogorov-Smirnoff statistics of 0.021, 0.018, and 0.017, and p-values of 0.75, 0.91, and 0.93, indicating that the null hypothesis is not rejected and the distributions are reasonably approximated by the Gaussian, .
@@
When the number of template libraries is small (as in the case considered here), it is prudent to adopt Bessel’s correction to get the unbiased estimators (see Appendix C for a derivation)
For a sample of galaxies or multiple spatial bins within a galaxy, in addition to measuring the average the mean , we can estimate the covariance matrix as
@@
For a small number of stellar libraries, again, we recommend Bessel’s corrected estimator, which is expressed as